# Semiparametric conformal prediction

### v1

@@ 5.1 Nonparametric density estimation
Motivated by recent developments in nonparametric estimation of copulas [42], we propose the use of nonparametric vine copulas described in Section 4 to model the distribution $F$ of score vectors. First, we model the marginals using the empirical CDF:
${\hat{F}}_{j}(s)=\frac{1}{n+1}\sum_{i\in\mathcal{I}_{\rm cal}\cup\{\infty\}}\mathbbm{1}[S_{j}^{(i)}\leq s],$ (12)
@@
For the bivariate copulas serving as building blocks for the vine, we opt for the transformation local likelihood kernel estimator (TLL), which allows KDE to be applied on the $[0,1]^{2}$ constrained copula space [53]. The resulting nonparametric vine copula ${\hat{C}}:[0,1]^{d}\to[0,1]$ yields the final estimated CDF
@@
Parameters of the copula density, including the kernel bandwidth and vine structure, are fit using the Akaike information criterion (AIC), a consistent model selection criterion for nonparametric models [54]. The CDF can then be obtained by pairwise Monte-Carlo (MC) integration of the density. We use the pyvinecopulib library [55] for fitting and model selection.
@@ 5.2 Semiparametric correction
$\Psi({\hat{F}})={\rm inf}\{s\in\mathbb{R}^{d}:{\hat{F}}(s)\geq 1-\alpha\}.$ (15)
This represents the $(1-\alpha)$-level curve of $\hat{F}$ illustrated in Figure 2. Note that the quantile is not unique; any quantile on the $(1-\alpha)$-level curve will do. The choice of which quantile to use for calibration depends on our preferences on the relative efficiency across the multiple targets. We can obtain a point on the level curve by running optimization to find the quantile in the copula space and mapping it back to the score space via the inverse PIT:
@@
We use the CMA-ES algorithm [56] to optimize the copula in Equation 16.
@@
$\psi_{\hat{C}}(\tilde{u})=\frac{(1-\alpha)-\mathbbm{1}[\tilde{u}\preccurlyeq U^{*}]}{|\nabla{\hat{C}}(U^{*})|^{2}}\nabla{\hat{C}}(U^{*}),$ (18)
where $\mathbbm{1}[b]$ is the indicator function that evaluates to 1 if $b$ is true and 0 otherwise. The derivation is provided in Section 10. The one-step estimator of $U^{*}$ is thus
@@
Figure 4 illustrates one-step estimation in the single-target case where we seek to estimate the $0.91-\alpha=$ 0.9 quantile (green dashed line) of the true score distribution (green curve) based on $n{\rm=}20$ samples. Standard CP takes the empirical CDF (black curve) and extracts the $\lceil(n+1)*(1-\alpha)\rceil/(n+1)$ quantile (dashed black line). Our approach is to “augment” the scores by fitting a nonparametric model for the CDF (purple curve). Here, the plug-in quantile (dashed purple line) lies very close to the empirical one and both overestimate the quantile, whereas the corrected estimate (solid purple line) offers a better approximation.
@@
The final prediction set for our method is given in Equation 6. As with [18], it takes hyperrectangular shapes. For $V(X,Y,\hat{f})_{j}=|Y_{j}-{\hat{f}}(X_{j})|$, for example, the target-wise intervals can be written
@@ 6 EXPERIMENTS
We investigate split conformal calibration schemes for the regression problem with $d$ real-valued targets. The underlying algorithm ${\hat{f}}$ fit on $\mathcal{I}_{\rm train}$ is a multi-task Lasso point predictor [57]. For calibration schemes that operate on vector scores, including ours, we choose the $L_{1}$ non-conformity score in Equation 4. Scores are evaluated on the same calibration set of size $n=|\mathcal{I}_{\rm cal}|$. We fix $0.91-\alpha=$ 0.9 and randomly split the data into training, calibration, and test sets.
@@
  • calibration applied to a scalar score defined as the $L_{2}$ norm [^3] This score definition yields prediction sets shaped as $d$-dimensional balls. The $L_{1}$ norm would yield cross-polytopes ($d$-dimensional generalization of diamonds). of the prediction error, $V(X,Y,\hat{f})=||Y-{\hat{f}}(X)||_{2}$ [29]
  • empirical copula fit on vector scores with the constraint that $U^{*}_{1}=\cdots=U^{*}_{d}$ [18]
@@
${\rm Coverage}=\frac{1}{|\mathcal{I}_{\rm test}|}\sum_{i\in\mathcal{I}_{\rm test}}\mathbbm{1}[Y^{(i)}\in{\Gamma}_{1-\alpha}(X^{(i)})].$ (21)
@@ 6.1 Synthetic example
We generate data using the penicillin production simulator from [58]. This function takes in seven input parameters controlling penicillin production and outputs three real-valued targets ($d{\rm=}3$), namely penicillin yield, production time, and carbon dioxide (CO2) emission. To simulate measurement errors, we add random Gaussian errors with scale equal to 10% of the noiseless function output, so that $Y=f_{0}(X)+\varepsilon$ where $f_{0}$ is the test function and $\varepsilon\sim\mathcal{N}(0,(f_{0}(X)/10)^{2})$. More details are in Section 13.1.
@@
On the other hand, the independent scheme is forced to extract an empirical marginal quantile at the $\lceil(0.9)^{1/d}(n+1)\rceil/(n+1)\approx 0.97$ level from only $n{\approx}100$ calibration instances, so results in overconservative prediction sets [18]. It cannot take advantage of the correlation in the residuals, particularly in the tails, that arise due to (tail) correlations in the target quantities, as visualized in Figure 8 and Figure 9.
@@ 6.2 Real datasets
Method Stock [59] ($d=6,n=63$) Caco2+ [60, 61] ($d=6,n=137$) rf1 [62] ($d=8,n=225$)
@@
We report performance on a variety of real-world datasets with 6, 8, or 16 targets and a range of calibration-set sizes in Table 2. In all of these problems, our corrected scheme is valid and has empirical coverage closest to 0.9. It is also the most efficient among the valid schemes. For Stock, Caco2+, and rf2, the correction is instrumental in making the plug-in estimate valid (above 0.9 coverage). At the same time, it does not significantly hurt efficiency.
@@ 7 LIMITATIONS
Just like the algorithm of [18], our algorithm does not formally guarantee finite-sample validity. Loss of coverage can occur if the estimated copula $\hat{C}$ of the scores, constructed with $n$ data points, deviates significantly from the “true” copula in the $n\to\infty$ limit. The parametric copula families considered in [18] is prone to model misspecification bias; we aim at minimizing this using nonparametric vines.
In our work, we used CMA-ES [56], a gradient-free evolutionary algorithm, to optimize the copula. This step can be made more efficient, particularly for large $d$, by using gradient-based algorithms. Unlike empirical copulas and some classes of parametric copulas, which lend themselves to gradient descent algorithms [22], nonparametric vines are more complicated for gradient computation. We leave this to future work.
@@ 8 CONCLUSION
An interesting avenue for future work may be to specify the structure of the vine based on domain knowledge about the target dependence structure. We may use C-vines for when one target variable dominates or D-vines for temporal variables considered in [22].
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Semiparametric Conformal Prediction:
@@ 9.1 Copulas
Using parametric marginal models for $F_{j},\forall j\in[d]$ amounts to an inference for margins (IFM) approach [63]. In this work, we use the empirical marginal distribution [64].
@@ 9.2 Vine copulas
Vine copulas are flexible models composed of bivariate building blocks. We will illustrate the pair copula decomposition for $d=3$, closely following the treatment in [65]. We can factorize the joint density as follows:
@@
If parametric copula families are insufficient for fitting the pseudo observations, nonparametric approaches offer a more flexible fit. Transformation local likelihood kernel estimators (TLL) perform well [53, 21]. We opt for the vine copula with TLL building blocks in this work.
The PCC described above generalizes to general $d$. The regular vine (R-vine) tree structure provides a framework for defining valid factorizations [66, 20]. A regular vine takes the form of a linked tree sequence, where the edges in one tree turn into the nodes of the next tree. Formally, a set of graphs $\mathcal{V}=(T_{1},\dotsc,T_{d-1})$ is a regular vine on $d$ elements if it satisfies the following:
@@
Theorem 2 (Existence of a regular vine distribution [20]).
@@
The parameters of the bivariate copulas can be estimated via maximum likelihood. While the number of valid vine structures grows super-exponentially with $d$, as $d!\times 2^{(d-2)(d-3)/2-1}$ [67], the greedy selection algorithm by [68] enables efficient exploration.
@@ 10.1 Univariate quantile
The derivation here closely follows the presentation in [25]. The estimand of interest is $\Psi_{1-\alpha}(P)=F^{-1}(1-\alpha)$. The quantile can be implicitly written
@@
With $F_{t}(s)=(1-t)F(s)+t\mathbbm{1}[{\tilde{s}}\leq s]$,
@@
$=\frac{F(F_{t}^{-1}(1-\alpha))-\mathbbm{1}[\tilde{s}\leq\Psi_{1-\alpha}(P_{t})]}{f_{t}\left(\Psi_{1-\alpha}(P_{t})\right)}.$
@@
$\psi_{P}({\tilde{s}})$ $=\frac{1-\alpha-\mathbbm{1}[\tilde{s}\leq\Psi_{1-\alpha}(P)]}{f\left(\Psi_{1-\alpha}(P)\right)}$
$=\frac{\mathbbm{1}[\Psi_{1-\alpha}(P)\leq{\tilde{s}}]-\alpha}{f\left(\Psi_{1-\alpha}(P)\right)}.\qed$
The pinball loss used for quantile regression is $l_{1-\alpha}(x)=x(\mathbbm{1}[x\geq 0]-\alpha)$ [69]. In the numerator of the last expression, one may recognize its derivative $l^{\prime}_{1-\alpha}(x)=\mathbbm{1}[x\geq 0]-\alpha$, namely $l^{\prime}_{1-\alpha}({\tilde{s}}-\Psi_{1-\alpha}(P))$.

### v2

@@ 5.1 Nonparametric density estimation
Motivated by recent developments in nonparametric estimation of copulas (Chen and Huang,, 2007), we propose the use of nonparametric vine copulas described in Section 4 to model the distribution $F$ of score vectors. First, we obtain the empirical marginal CDF:
${\hat{F}}_{j}(s)=\frac{1}{n+1}\sum_{i\in\mathcal{I}_{\rm cal}\cup\{\infty\}}\mathds{1}[S_{j}^{(i)}\leq s],$ (12)
@@
For the bivariate copulas serving as building blocks for the vine, we opt for the transformation local likelihood kernel estimator (TLL), which allows KDE to be applied on the $[0,1]^{2}$ constrained copula space (Geenens,, 2014). The resulting nonparametric vine copula ${\hat{C}}:[0,1]^{d}\to[0,1]$ yields the final estimated CDF
@@
Parameters of the copula density, including the kernel bandwidth and vine structure, are fit using the Akaike information criterion (AIC), a consistent model selection criterion for nonparametric models (Akaike,, 1974). The CDF can then be obtained by pairwise Monte-Carlo (MC) integration of the density. We use the pyvinecopulib library (Vatter and Nagler,, 2024) for fitting and model selection.
@@ 5.2 Semiparametric correction
$\Psi({\hat{F}})=\{s\in\mathbb{R}^{d}:{\hat{F}}(s)\geq 1-\alpha\}.$ (15)
This represents the $(1-\alpha)$-level curve of $\hat{F}$ illustrated in Figure 2. Note that the quantile is not unique; any quantile on the $(1-\alpha)$-level curve will do. The choice of which quantile to use for calibration depends on our preferences on the relative efficiency across the multiple targets. Suppose we would like a quantile with the smallest $L_{1}$ norm. We can obtain this point on the level curve by running optimization to find the quantile in the copula space and mapping it back to the score space via the inverse PIT:
@@
We use the CMA-ES algorithm (Hansen et al.,, 2003) to optimize the copula in Equation 16.
@@
$\psi_{\hat{C}}(\tilde{u})=\frac{(1-\alpha)-\mathds{1}[\tilde{u}\preccurlyeq U^{*}]}{|\nabla{\hat{C}}(U^{*})|^{2}}\nabla{\hat{C}}(U^{*}),$ (18)
where $\mathds{1}[b]$ is the indicator function that evaluates to 1 if $b$ is true and 0 otherwise. The derivation is provided in Section 10. The one-step estimator of $U^{*}$ is thus
@@
Figure 4 illustrates one-step estimation in the single-target case where we seek to estimate the $0.91-\alpha=$ 0.9 quantile (green dashed line) of the true score distribution (green curve) based on $n{\rm=}20$ samples. Standard CP takes the empirical CDF (black curve) and extracts the $\lceil(n+1)(1-\alpha)\rceil/(n+1)$ quantile (dashed black line). Our approach is to “augment” the scores by fitting a nonparametric model for the CDF (purple curve). Here, the plug-in quantile (dashed purple line) lies very close to the empirical one and both overestimate the quantile, whereas the corrected estimate (solid purple line) offers a better approximation.
@@
The final prediction set for our method is given in Equation 6. As with Messoudi et al., (2021), it takes hyperrectangular shapes. For $V(X,Y,\hat{f})_{j}=|Y_{j}-{\hat{f}}(X_{j})|$, for example, the target-wise intervals are
@@
5.3 Theoretical results
We informally state the theoretical guarantees of our method below. The coverage of Algorithm 1 is asymptotically exact by Theorem 2 and is approximately valid by Theorem 3. Complete proofs can be found in Section 13. We assume the following:
  • The true distribution $F^{*}$ over the score vectors $S$ is absolutely continuous.
[…]
Theorem 2 (Asymptotic consistency).
Under the stated regularity conditions, prediction set $\Gamma_{1-\alpha}$ from Algorithm 1 constructed using the one-step correction $U^{*}$ satisfies
$\mathbb{P}[Y^{*}\in\Gamma_{1-\alpha}(X^{*})\to 1-\alpha\textrm{ as }n\to\infty,$
for a test point $\left(X^{*},Y^{*}\right)$. That is, its coverage is asymptotically exact.
The statement follows from the uniform consistency of the copula estimator $\hat{F}$, which implies consistency of its quantile, and asymptotic normality of the one-step estimator.
Theorem 3 (Approximate validity).
Suppose that the total variation (TV) distance between the true distribution $F^{*}$ and the estimated distribution $\hat{F}$ is bounded by $\epsilon$. That is,
$\sup_{S}|F^{*}(S)-{\hat{F}}(S)|\leq\epsilon.$
Then the resulting prediction set $\Gamma_{1-\alpha}$ from Algorithm 1 constructed either from the plug-in quantile $U$ or the one-step corrected quantile $U^{*}$ is marginally valid at the $1-\alpha-\epsilon$ level:
$\mathbb{P}[Y^{*}\in\Gamma_{1-\alpha}(X^{*})]\geq 1-\alpha-\epsilon.$
The proof proceeds as follows. By the data processing inequality, the TV distance between $F^{*}$ and $\hat{F}$ upper-bounds the TV distance between their corresponding quantiles. The same holds between the quantile of $\hat{F}$ and the one-step corrected quantile, as the correction term does not add additional information.
Finite-sample validity.
In Section 12, we also introduce Algorithm 2, a variant of Algorithm 1 that further splits the calibration set into $\mathcal{I}_{{\rm cal},1}$ and $\mathcal{I}_{{\rm cal},2}$. The set $\mathcal{I}_{{\rm cal},1}$ is used to build the empirical CDF $\hat{F}_{j},\forall j\in[d]$ and the set $\mathcal{I}_{{\rm cal},2}$ with $n^{\prime}=|\mathcal{I}_{{\rm cal},2}|$ is used to fit the empirical copula $\hat{C}$, from which the $\lceil(1-\alpha)(n^{\prime}+1)/(n^{\prime}+1)\rceil$ quantile is extracted. The extra split provides finite-sample validity by Theorem 5, because it makes the vector marginals $U$ between $\mathcal{I}_{{\rm cal},2}$ and the test set exchangeable.
@@ 6 EXPERIMENTS
We investigate split conformal calibration schemes for regression with $d$ real-valued targets. The underlying algorithm ${\hat{f}}$ fit on $\mathcal{I}_{\rm train}$ is a multi-task Lasso point predictor (Tibshirani,, 1996). For calibration schemes that operate on vector scores, including ours, we choose the $L_{1}$ non-conformity score in Equation 4. Scores are evaluated on the same calibration set of size $n=|\mathcal{I}_{\rm cal}|$. We fix $0.91-\alpha=$ 0.9 and randomly split the data into training, calibration, and test sets.
@@
  • calibration applied to a scalar score defined as the $L_{2}$ norm [^3] This score definition yields prediction sets shaped as $d$-dimensional balls. The $L_{1}$ norm would yield cross-polytopes ($d$-dimensional generalization of diamonds). The $L_{\infty}$ norm would yield hypercubes. of the prediction error, $V(X,Y,\hat{f})=||Y-{\hat{f}}(X)||_{2}$ (Feldman et al.,, 2023)
  • empirical copula fit on vector scores with the constraint $U^{*}_{1}=\cdots=U^{*}_{d}$ Messoudi et al., (2021)
@@
${\rm Coverage}=\frac{1}{|\mathcal{I}_{\rm test}|}\sum_{i\in\mathcal{I}_{\rm test}}\mathds{1}[Y^{(i)}\in{\Gamma}_{1-\alpha}(X^{(i)})].$ (21)
@@ 6.1 Synthetic example
We generate data using the penicillin production simulator from (Liang and Lai,, 2021). This function takes in seven input parameters controlling penicillin production and outputs three real-valued targets ($d{\rm=}3$), namely penicillin yield, production time, and carbon dioxide emission. To simulate measurement errors, we add random Gaussian errors with scale equal to 10% of the noiseless function output, so that $Y=f_{0}(X)+\varepsilon$ where $f_{0}$ is the test function and $\varepsilon\sim\mathcal{N}(0,(f_{0}(X)/10)^{2})$. More details are in Section 15.1.
@@
On the other hand, the independent scheme is forced to extract an empirical marginal quantile at the $\lceil(0.9)^{1/d}(n+1)\rceil/(n+1)\approx 0.97$ level from only $n{\approx}100$ calibration instances, so become overconservative (Messoudi et al.,, 2021). It cannot take advantage of the correlation in the residuals, particularly in the tails, that arise due to (tail) correlations in the target quantities, as visualized in Figure 8 and Figure 9.
@@
Section 14 includes a complete set of experiments including our Algorithm 2 as well as other types of scalar scores ($L_{2}$ and $L_{\infty}$ norms) and a “split” scalar-score method that further splits the calibration set for normalizing the scores prior to scalarization.
@@
Plug-in, split (ours) $0.92\pm 0.02$ $3.6\pm 0.3$
Corrected, split (ours) $0.90\pm 0.03$ $3.5\pm 0.4$
@@ 6.2 Real datasets
Method Stock ($d=6,n=63$) Caco2+ ($d=6,n=137$) rf1 ($d=8,n=225$)
@@
We report performance on a variety of real-world datasets with 6, 8, or 16 targets and a range of calibration-set sizes in Table 2. In all of these problems, our corrected scheme is valid and has empirical coverage closest to 0.9. It is also the most efficient among the valid schemes. For Stock (Liu and Yeh,, 2017), Caco2+ (Wang et al.,, 2016; Park et al.,, 2024), and rf2 (Spyromitros-Xioufis et al.,, 2016), the correction is instrumental in making the plug-in estimate valid (above 0.9 coverage). At the same time, it does not significantly hurt efficiency.
@@ 7 LIMITATIONS
Just like the algorithm of Messoudi et al., (2021), our algorithm only guarantees approximate finite-sample validity. Loss of coverage can occur if the estimated copula $\hat{C}$ of the scores, constructed with $n$ data points, deviates significantly from the “true” copula in the $n\to\infty$ limit. The parametric copula families considered in Messoudi et al., (2021) is prone to model misspecification bias; we aim at minimizing this using nonparametric vines.
In our work, we used CMA-ES (Hansen et al.,, 2003), a gradient-free evolutionary algorithm, to optimize the copula. This step can be made more efficient, particularly for large $d$, by using gradient-based algorithms. Unlike empirical copulas and some classes of parametric copulas, which lend themselves to gradient descent algorithms (Sun and Yu,, 2024), nonparametric vines are more complicated for gradient computation. We leave this to future work.
@@ 8 CONCLUSION
An interesting avenue for future work may be to specify the structure of the vine based on domain knowledge about the target dependence structure. We may use C-vines for when one target variable dominates or D-vines for temporal variables considered in (Sun and Yu,, 2024).
Acknowledgements
We thank Anastasios Angelopoulos for the insightful feedback.
@@ 9.1 Copulas
Using parametric marginal models for $F_{j},\forall j\in[d]$ amounts to an inference for margins (IFM) approach (Joe,, 2005). In this work, we use the empirical marginal distribution (Genest et al.,, 1995).
@@ 9.2 Vine copulas
Vine copulas are flexible models composed of bivariate building blocks. We will illustrate the pair copula decomposition for $d=3$, closely following the treatment in Czado and Nagler, (2022). We can factorize the joint density as follows:
@@
If parametric copula families are insufficient for fitting the pseudo observations, nonparametric approaches offer a more flexible fit. Transformation local likelihood kernel estimators (TLL) perform well (Geenens,, 2014; Nagler et al.,, 2017). We opt for the vine copula with TLL building blocks in this work.
The PCC described above generalizes to general $d$. The regular vine (R-vine) tree structure provides a framework for defining valid factorizations (Bedford and Cooke,, 2001, 2002). A regular vine takes the form of a linked tree sequence, where the edges in one tree turn into the nodes of the next tree. Formally, a set of graphs $\mathcal{V}=(T_{1},\dotsc,T_{d-1})$ is a regular vine on $d$ elements if it satisfies the following:
@@
Theorem 4 (Existence of a regular vine distribution (Bedford and Cooke,, 2002)).
@@
The parameters of the bivariate copulas can be estimated via maximum likelihood. While the number of valid vine structures grows super-exponentially with $d$, as $d!\times 2^{(d-2)(d-3)/2-1}$ (Napoles et al.,, 2010), the greedy selection algorithm by Dissmann et al., (2013) enables efficient exploration.
@@ 10.1 Univariate quantile
The derivation here closely follows the presentation in Hines et al., (2022). The estimand of interest is $\Psi_{1-\alpha}(P)=F^{-1}(1-\alpha)$. The quantile can be implicitly written
@@
With $F_{t}(s)=(1-t)F(s)+t\mathds{1}[{\tilde{s}}\leq s]$,
@@
$=\frac{F(F_{t}^{-1}(1-\alpha))-\mathds{1}[\tilde{s}\leq\Psi_{1-\alpha}(P_{t})]}{f_{t}\left(\Psi_{1-\alpha}(P_{t})\right)}.$
@@
$\psi_{P}({\tilde{s}})$ $=\frac{1-\alpha-\mathds{1}[\tilde{s}\leq\Psi_{1-\alpha}(P)]}{f\left(\Psi_{1-\alpha}(P)\right)}$
$=\frac{\mathds{1}[\Psi_{1-\alpha}(P)\leq{\tilde{s}}]-\alpha}{f\left(\Psi_{1-\alpha}(P)\right)}.\qed$
The pinball loss used for quantile regression is $l_{1-\alpha}(x)=x(\mathds{1}[x\geq 0]-\alpha)$ (Koenker and Bassett Jr,, 1978). In the numerator of the last expression, one may recognize its derivative $l^{\prime}_{1-\alpha}(x)=\mathds{1}[x\geq 0]-\alpha$, namely $l^{\prime}_{1-\alpha}({\tilde{s}}-\Psi_{1-\alpha}(P))$.